OpenStudy (anonymous):

Use the simple interest formula to find the missing value. Rates are annual rates. Assume 360 days per year and 30 days per month. p = ?, r = 4.5%, t = 10 days, i = $26.68 A. p = $21,344.00 B. p = $711.47 C. p = $7114.67 D. p = $21,670.09

OpenStudy (phi):

the simple interest formula is I = P*r*t we need to replace the letters with the numbers given in the problem I= 26.68 r= 4.5% annually. We must change this to a decimal by dividing by 100 r= 0.045 t is time. It must be in years. the problem says t=10 days, in years that is 10 days/ 360 days or 1/36 26.68= P*0.045*1/36 to find P, multiply both sides by 36/0.045 26.68*36/0.045 = P type 26.68*36/0.045 = in the google search window, or use a calculator to find P

OpenStudy (anonymous):

thanks

OpenStudy (phi):

Do you think you could do another problem like this?

OpenStudy (anonymous):

no i not good at math